OpenStudy (anonymous):

find the x-coordinate of all points on the curve y=36xsin(17x) + 306sqrt3(x^2) -33 domain pi/34 < x < 3pi/34 where the tangent line passes through the point (0, -33)

OpenStudy (anonymous):

\[y = 36x \sin(17x) + 306\sqrt{3}x^2 -33, \frac{ \Pi }{ 34 } < x < \frac{ 3\Pi }{ 34 }\]

OpenStudy (anonymous):

im getting this for the derivative \[y' = 612xcos(17x) + 36\sin(17x) + 712\sqrt{3}x \]